OpenStudy (r_josh10):

If y varies directly with x, and y = 6 when x = 9, what is the value of x when y = 7 I need Help

OpenStudy (anonymous):

y = kx Plug in what we know: 6 = 9k Divide 9 to both sides: 6/9 = k k = 2/3 Okay, now let's plug in our constant with y = 7. y = kx 7 = 2/3x Divide 2/3 to both sides or multiply by its reciprocal, 3/2. So what's 3/2 * 7?

OpenStudy (r_josh10):

10.5?

OpenStudy (anonymous):

Yes, that's our answer. x is 10.5 when y is 7.

OpenStudy (r_josh10):

Ok thanks @iGreen.